OpenStudy (anonymous):

the half-life of plutonium-239 is 24,100 years what is its initial amount if after 1000 years its amount is 2.1 grams?

OpenStudy (anonymous):

2.4g I'll explain in a minute.

OpenStudy (anonymous):

Sorry, 2.04g

OpenStudy (anonymous):

This is exponential decay. The rate of growth or decay in something undergoing this kind of change is in proportion to itself. That is,\[\frac{dN}{dt} =n^2N\]where I've written n^2 as the constant of proportionality to emphasize it's magnitude (we can worry about signs later). You solve this type of equation as\[\frac{dN}{N}=n^2dt \rightarrow \int\limits_{}^{}\frac{dN}{N}=\int\limits_{}{}n^2dt \rightarrow \ln N=n^2t + c\]

OpenStudy (anonymous):

Exponentiate both sides to get\[N=e^{n^2t+c}=e^ce^{n^2t}=Ce^{n^2t}\]

OpenStudy (anonymous):

When t=0, the equation is N(0)=C, so C is the initial amount. Also, when t=24100 years, half of the initial substance is gone, so at this time,\[\frac{1}{2}N(0)=N(0)e^{n^2(24,100)} \rightarrow \frac{1}{2}=e^{n^2(24,100)}\]take the natural log of both sides to solve for n^2:\[\ln \frac{1}{2}=n^2(24100)yr \rightarrow n^2=\frac{\ln 1/2}{24100yr}=-\frac{\ln 2}{24000}\]

OpenStudy (anonymous):

So your equation for the amount of plutonium-239 at time t is:\[N(t)=2.1 e^{-\frac{\ln 2}{24100yr}t} grams\]where t is in years.

OpenStudy (anonymous):

For t=1000 years,\[N(1000)=2.1 e^{-\ln 2/24100yr \times 1000yr}g \approx 2.04g\]

OpenStudy (anonymous):

Okay?

OpenStudy (anonymous):

There's a typo. above -> 24000 should be 24100.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Is this your first time here?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

We like to be thanked in fan points ;) There should be a link next to my name, "Become a fan".